a car travelling at 22.5m/s decelerate at 2.27m/s². how far does it go in third minutes after the brakes were applied

1
Expert's answer
2022-05-24T13:56:17-0400
"s(t)=v_0t-at^2\/2=22.5t-1.135t^2"


"d=s(3)-s(2)\\
=(22.5*3-1.135*3^2)-(22.5*2-1.135*2^2)\\
=16.8\:\rm m"